OpenStudy (sissyedgar):

Help! ;) Solve. 4y + 16 > –8 A. y > –6 B. y > –2 C. y > 2 D. y > 6

OpenStudy (sissyedgar):

@dan815

OpenStudy (misty1212):

HI!!

OpenStudy (sissyedgar):

hey

OpenStudy (misty1212):

subtract \(16\) first

OpenStudy (misty1212):

\[4y>-8-16\\ 4y>-24\]

OpenStudy (misty1212):

then divide by 4 and you are done!

OpenStudy (sissyedgar):

-6?

OpenStudy (misty1212):

yeah

OpenStudy (misty1212):

well \(y>-6\)

OpenStudy (sissyedgar):

ty!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (sissyedgar):

can u help wit another? ♥

OpenStudy (misty1212):

sure!!

OpenStudy (misty1212):

you want to just post it here?

OpenStudy (sissyedgar):

mhm

OpenStudy (misty1212):

its gotta be one of the first ones, a "greater than" since it is the stuff to the right

OpenStudy (misty1212):

it is hard for me to see if that little circle is filled in or is open

OpenStudy (sissyedgar):

o, ok.. makes since. thx

OpenStudy (misty1212):

if it is filled in, use \(\geq\) and if it is open use \(>\)

OpenStudy (misty1212):

looks like it is probably open, so go with A

OpenStudy (sissyedgar):

ok

OpenStudy (sissyedgar):

2 more!

OpenStudy (misty1212):

lol lets do them!!

OpenStudy (misty1212):

\[\huge -\frac{t}{5}\geq15\] multiply by \(-5\) and change the inequality

OpenStudy (misty1212):

\[\huge t\leq -75\]

OpenStudy (sissyedgar):

ok. ty so much! last 1! YAY!

OpenStudy (sissyedgar):

When its over ill tell u my score :)

OpenStudy (misty1212):

oooh this one you gotta solve 4 inequalities in one problem!!

OpenStudy (sissyedgar):

ik

OpenStudy (sissyedgar):

well, i guess 4 more then.. not 1! :P

OpenStudy (misty1212):

ok lets knock em out \[2x+8>6\\ 2x>-2\\x>-1\] do you see that one ?

OpenStudy (sissyedgar):

A.?

OpenStudy (sissyedgar):

A.C.D. are a lot alike

OpenStudy (sissyedgar):

well.. all of em

OpenStudy (misty1212):

nah it is greater than so goes to the right plus it is \(>\) so the circle is open

OpenStudy (sissyedgar):

D.

OpenStudy (misty1212):

i can't really read the numbers but it looks iike D

OpenStudy (misty1212):

ok on to the next one

OpenStudy (sissyedgar):

ok

OpenStudy (sissyedgar):

want me 2 resend the file with bigger numbers?

OpenStudy (misty1212):

\[6<4-2x\\ 2<-2x\\ -1>x\]

OpenStudy (misty1212):

nah its ok

OpenStudy (sissyedgar):

okie dokie

OpenStudy (misty1212):

for this one pick the arrow that goes to the left

OpenStudy (misty1212):

and circle is open

OpenStudy (sissyedgar):

A.

OpenStudy (misty1212):

kk A

OpenStudy (misty1212):

next

OpenStudy (misty1212):

\[9\leq -11x-2\\ 11\leq -11x\\ -1\geq x\]

OpenStudy (sissyedgar):

C.?

OpenStudy (misty1212):

this one goes to the left, circle is closed in

OpenStudy (misty1212):

nah B

OpenStudy (sissyedgar):

so B.?

OpenStudy (misty1212):

goes to the left

OpenStudy (sissyedgar):

ok, well then the last 1 has 2 b C.!

OpenStudy (misty1212):

last one is whatever is left, unless they are trying to trick you

OpenStudy (sissyedgar):

lol,, ok lemme c wat my score is!

OpenStudy (misty1212):

yeah is C

OpenStudy (sissyedgar):

100%!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (sissyedgar):

Thankies!!!

OpenStudy (sissyedgar):

are u gonne b on a lil while longer?

OpenStudy (misty1212):

lol how much more you got?

OpenStudy (sissyedgar):

if so i need help wit another one in a few mins... brb i gotta P! :P

OpenStudy (misty1212):

kk i wait

OpenStudy (sissyedgar):

bk... gimme like 3 more mins tho! ok?

OpenStudy (misty1212):

k

OpenStudy (sissyedgar):

ok, its a different thing so imma do it in a new question! :)

OpenStudy (sissyedgar):

its 12 questions!